### v4

# Multilingual Machine Translation with Open Large Language Models at Practical Scale: An Empirical Study

Abstract
Large language models (LLMs) have shown continuously improving multilingual capabilities, and even small-scale open-source models have demonstrated rapid performance enhancement. In this paper, we systematically explore the abilities of open LLMs with less than ten billion parameters to handle multilingual machine translation (MT) tasks. We conduct comprehensive evaluations on six popular LLMs and find that models like Gemma2-9B exhibit impressive multilingual translation capabilities. We then introduce the Parallel-First Monolingual-Second (PFMS) data mixing strategy in the continual pretraining stage to further enhance the MT performance and present GemmaX2-28, a 9B model achieving top-tier multilingual translation performance across 28 languages111Arabic (ar), Bengali (bn), Czech (cs), German (de), English (en), Spanish (es), Persian (fa), French (fr), Hebrew (he), Hindi (hi), Indonesian (id), Italian (it), Japanese (ja), Khmer (km), Korean (ko), Lao (lo), Malay (ms), Burmese (my), Dutch (nl), Polish (pl), Portuguese (pt), Russian (ru), Thai (th), Tagalog (tl), Turkish (tr), Urdu (ur), Vietnamese (vi), and Chinese (zh).. Specifically, GemmaX2-28 consistently outperforms the state-of-the-art (SOTA) models such as TowerInstruct Alves et al. (2024) and X-ALMA Xu et al. (2024b) and achieves competitive performance with Google Translate and GPT-4-turbo. 222Models are released at https://huggingface/GemmaX2.
Multilingual Machine Translation with Open Large Language Models at Practical Scale: An Empirical Study
Menglong Cui††thanks: Equal contribution. Correspondence to: Pengzhi Gao <gaopengzhi@xiaomi.com>., Pengzhi Gao11footnotemark: 1, Wei Liu, Jian Luan, Bin Wang Xiaomi Inc., Beijing, China cuimenglongcs@gmail.com {gaopengzhi,liuwei40,luanjian,wangbin11}@xiaomi.com
1 Introduction
Large language models (LLMs) such as GPT models OpenAI (2023a, b), PaLM models Chowdhery et al. (2022); Anil et al. (2023), and others have demonstrated remarkable capabilities in multilingual translation Jiao et al. (2023); Vilar et al. (2023). Recently, significant efforts have been directed towards bolstering the multilingual competencies of open LLMs across a diverse linguistic spectrum, beyond English and Chinese. For example, over 5% of the LLaMA3 Dubey et al. (2024) pretraining datasets consist of high-quality non-English data covering over languages; Qwen2/2.5 Yang et al. (2024) claim to have multilingual support for over languages; Üstün et al. (2024) introduce Aya-101, a massively multilingual generative language model supporting languages.
Zhu et al. (2024b) evaluate popular LLMs on multilingual MT tasks and find that the translation capability of LLMs is continually evolving. However, the models assessed (XGLM Lin et al. (2022), OPT Zhang et al. (2022), Falcon Almazrouei et al. (2023), BLOOMZ Workshop et al. (2023), LLaMA2 Touvron et al. (2023b)) are outdated due to the rapid development of open LLMs, and the multilingual translation ability of the latest models remains unclear. One natural question arises: 1) How do the latest open LLMs with practical scale perform multilingual MT tasks?
To further boost LLMs’ translation capability, existing approaches usually adopt multilingual corpora during continual pretraining on open LLMs. Specifically, Xu et al. (2024a) utilize continual pretraining on monolingual datasets. Guo et al. (2024) apply continual pretraining on monolingual datasets followed by parallel datasets. Alves et al. (2024) perform continual pretraining on a multilingual mixture of monolingual (two-thirds) and parallel (one-third) datasets. However, the optimal approach to leverage monolingual and parallel datasets for multilingual MT remains under-explored. Another natural question arises: 2) What are the best practices for leveraging monolingual and parallel corpora to enhance the multilingual MT performance of LLMs?
In this paper, we first evaluate the in-context capabilities of the latest open-source LLMs including Mistral-7B-v0.3 Jiang et al. (2023), Qwen2/2.5-7B Yang et al. (2024), LLaMA3/3.1-8B Dubey et al. (2024), and Gemma2-9B Team et al. (2024) across languages on the multilingual MT tasks. We find that open LLMs like Gemma2-9B exhibit remarkable multilingual translation capabilities but still fall behind strong closed-source models. We then systematically investigate the most effective methods to leverage monolingual and parallel data for further boosting Gemma2-9B’s capabilities and propose the Parallel-First Monolingual-Second (PFMS) data mixing strategy. Leveraging PFMS strategy when continually pretraining on Gemma2-9B followed by instruction finetuning on a small set of high-quality translation pairs, we learn GemmaX2-28-9B, a many-to-many multilingual MT model that performs competitively with Google Translate and GPT-4-turbo across widely spoken languages. The contributions of this paper can be summarized as follows:
We benchmark the latest open-source LLMs on multilingual MT tasks in widely spoken languages, covering English-centric and Chinese-centric translation directions.
We systematically explore the optimal training recipe for boosting multilingual MT with LLMs and introduce the Parallel-First Monolingual-Second (PFMS) data mixing strategy for the continual pretraining stage.
We publicly release GemmaX2-28-9B, the many-to-many multilingual translation model supporting languages, which consistently outperforms open-source alternates and is competitive with closed-source models such as Google Translate and GPT-4-turbo.
2 Related Work
2.1 Multilinguality in LLMs
Large language models have gained substantial attention from the research community due to their outstanding performance on various tasks Winata et al. (2021); Wei et al. (2022); Touvron et al. (2023a). However, most advancements in LLMs have primarily focused on English or Chinese, resulting in inadequate performance in low-resource languages Ebrahimi et al. (2022); Asai et al. (2024). The reason is that the language distribution of the training data for LLMs is highly imbalanced and the quality varies across languages Ding et al. (2024). Numerous efforts have been made to enhance the multilingual capabilities of LLMs. Some studies seek to enhance the performance of LLMs on low-resource languages through continual pretraining Cui et al. (2024b) or supervised finetuning Üstün et al. (2024) using data from these languages. Additionally, some researchers Li et al. (2024a) leverage contrastive learning to align the internal representations of different languages, allowing the model to improve its cross-lingual capabilities with minimal training data. In our study, we assessed the latest open-source LLMs on translation tasks across languages. We find that these latest models still show significant performance disparities across different languages, indicating that low-resource languages continue to pose a significant challenge for these models. Therefore, we expand the Gemma2 model to cover commonly used languages, enabling it to demonstrate strong generative and translation capabilities across these languages.
2.2 Multilingual MT with LLMs
The use of LLMs has shown significant progress in MT tasks Lu et al. (2024); Xu et al. (2024c); Li et al. (2024b); Gao et al. (2024), leading to different approaches on LLM-based translation. One line of work focuses on the in-context translation capabilities, and LLMs are provided with parallel sentences to guide the model in generating the target sentence. Some studies Agrawal et al. (2023); Zhu et al. (2024a); Cui et al. (2024a) show that MT performance can be improved by using semantically related parallel sentences as examples, exhibiting promising results in scenarios with limited computational resources and insufficient parallel data.
Another line involves finetuning with translation instructions. Xu et al. (2024a) initially pretrained the base model on monolingual data, followed by finetuning on small human-written parallel datasets, which resulted in strong translation performance. Guo et al. (2024) challenge the perspective that the impact of parallel data is reduced in the LLM era and demonstrate the effectiveness of parallel data during continual pretraining on enhancing multilingual translation. However, the distribution of the parallel datasets could be highly imbalanced across different languages for massively multilingual translation scenarios. Alves et al. (2024) incorporates a fixed multilingual mixture of monolingual (two-thirds) and parallel (one-third) data, further enhancing LLMs’ translation capabilities. Our work differs in that we primarily focus on exploring the optimal mixing strategy of monolingual and parallel data during continual pretraining to achieve the best translation performance.
3 Datasets and Baseline Settings
In this section, we describe the evaluation datasets used in Sections 4 and 5 as well as the model configurations.
3.1 Datasets
We conduct experiments on languages across a broad linguistic spectrum. The detailed information of all languages is summarized in Table 3. We evaluate the multilingual translation performance on the FLORES-200 Team et al. (2022) benchmark. To avoid data leakage issues, we also consider the WMT-24 test sets for evaluating multilingual translation capabilities. Only sources are provided in the WMT-24 benchmark, and we adopt English sentences for reference-free evaluation.
3.2 Models
We evaluate the translation performance of six open-source LLMs: Mistral-7B-v0.3 Jiang et al. (2023), Qwen2/2.5-7B Yang et al. (2024), LLaMA3/3.1-8B Dubey et al. (2024) and Gemma2-9B Team et al. (2024). We also summarize the results of several SOTA models as follows:
Google Translate: The translation performance by leveraging the Google Translate API333https://translate.google.com/. We include it to represent the commercial MT system.
GPT-3.5/4 Turbo: The translation performance by leveraging the OpenAI API444https://api.openai.com/v1/chat/completions with greedy decoding for inference. Notably, we adopt the same in-context learning strategy as that employed by open-source LLMs.
NLLB-54.5B: The largest and best multilingual NMT model with encoder-decoder architecture released by the No Language Left Behind (NLLB) project Team et al. (2022).
3.3 Evaluation
For the FLORES-200 benchmark, we evaluate MT performance by spBLEU555We calculate the spBLEU scores via sacreBLEU Post (2018) with the flores200 tokenizer. Goyal et al. (2022) and COMET666https://huggingface.co/Unbabel/wmt22-comet-da Rei et al. (2020). For the WMT-24 benchmark, we adopt two reference-free models, XCOMET777https://huggingface.co/Unbabel/XCOMET-XXL Guerreiro et al. (2023) and COMETKiwi888https://huggingface.co/Unbabel/wmt23-cometkiwi-da-xxl Rei et al. (2023), each of which has B parameters and demonstrates high correlation with human judgments Freitag et al. (2023).
Model | WMT-24 | FLORES-200 | |||
en xx | en xx | xx en | zh xx | xx zh | |
Google Translate | 77.64 / 73.00 | 41.52 / 88.51 | 45.35 / 88.83 | 29.64 / 85.69 | 34.86 / 87.20 |
GPT-4-turbo | 77.55 / 70.68 | 37.01 / 87.14 | 41.53 / 88.40 | 24.66 / 84.31 | 28.80 / 85.90 |
GPT-3.5-turbo | 70.63 / 62.64 | 32.24 / 82.49 | 36.18 / 85.63 | 20.35 / 79.34 | 24.12 / 82.61 |
NLLB-54.5B | - | 37.34 / 87.05 | 43.63 / 88.32 | 25.41 / 84.42 | 20.73 / 80.72 |
Gemma2-9B | 72.06 / 67.05 | 33.05 / 84.65 | 42.00 / 87.94 | 20.54 / 80.54 | 27.77 / 85.34 |
Llama3.1-8B | 65.23 / 58.89 | 28.52 / 81.60 | 37.72 / 86.74 | 17.66 / 78.09 | 25.53 / 83.49 |
Llama3-8B | 65.02 / 58.10 | 27.93 / 81.28 | 37.39 / 86.59 | 16.83 / 77.40 | 22.81 / 83.11 |
Qwen2.5-7B | 59.36 / 51.59 | 22.65 / 75.20 | 34.64 / 84.94 | 14.97 / 72.64 | 25.30 / 83.16 |
Qwen2-7B | 58.12 / 49.68 | 21.69 / 75.68 | 34.22 / 85.15 | 14.44 / 73.19 | 25.45 / 83.55 |
Mistral-7B-v0.3 | 51.21 / 40.48 | 17.04 / 67.61 | 29.44 / 81.35 | 8.18 / 62.31 | 12.19 / 72.62 |
4 Benchmarking Open LLMs for Multilingual Machine Translation
In this section, we report empirical results on tokenizer efficiency and multilingual MT performance across various open-source LLMs.
4.1 Tokenizer Efficiency
The open-source LLMs are typically pretrained on one or a few dominant languages, and tokenization on low-resource languages usually results in undesirable long sequences of subwords which leads to excessive GPU memory consumption during training and slow decoding speed during inference.
Following the line of Liao et al. (2024), we compare the length differences between tokenized English sentences and their corresponding non-English counterparts to evaluate the tokenizer efficiency across different LLMs. Specifically, we define the length ratio as follows:
where represents the English sentence, and denotes the corresponding non-English sentence. A smaller length ratio is preferred, meaning that the tokenizer encodes the non-English sentences as efficiently as the English sentence.
We conduct our experiments on the FLORES-200 devtest dataset, which includes sentences for each language. Note that LLaMA3 and LLaMA3.1 share the same tokenizer, as do Qwen2 and Qwen2.5. We also include the tokenizer of the NLLB-54.5B model as a strong baseline for comparison. The averaged length ratio for each language is illustrated in Figure 1. By checking the length ratio for various languages, we have the following observations: 1) Due to its extensive support and optimization for more than languages, NLLB-54.5B has the most balanced tokenizer and achieves low length ratios across different languages. 2) Open-source LLMs demonstrate notably high length ratios for low-resource languages such as Khmer, Lao, and Burmese. Gemma2-9B consistently achieves better length ratios for almost all languages compared with other LLMs.
4.2 In-context Multilingual Translation Performance with Open LLMs
We here investigate the multilingual translation performance of different LLMs with in-context learning strategies on the FLORES-200 and WMT-24 benchmarks. For each open LLM, we report its translation performance across languages with five randomly selected translation pairs999We apply the same five randomly selected translation pairs as exemplars for each direction during evaluation. from the FLORES-200 development dataset as the in-context exemplars. Note that we also adopt the FLORES-200 development dataset for the WMT-24 benchmark. Following the line of Zhu et al. (2024b), we adopt the format “<X>=<Y>” as the in-context template, where <X> and <Y> denote the source and target sentences of the select parallel sentence pairs. All experiments are conducted based on OpenICL101010https://github.com/Shark-NLP/OpenICL Wu et al. (2023).
We report the averaged multilingual translation performance on the WMT-24 and FLORES-200 benchmarks in Tables 1. By checking the translation performance for various languages and models, we have the following observations:
Open LLMs demonstrate impressive multilingual translation capability. Specifically, Gemma2-9B performs the best among the open LLMs evaluated and outperforms GPT-3.5-turbo on average.
Open LLMs still lag behind the strong supervised NMT models such as NLLB-54.5B, especially in low-resource languages.
Google Translate exhibits remarkable multilingual translation performance and even outperforms GPT-4-turbo on both the FLORES-200 and WMT-24 benchmarks.
We do not observe serious data leakage issues on public datasets (FLORES-200) for the open LLMs we evaluated, where the translation performance on both WMT-24 and FLORES-200 benchmarks share a similar trend.
We then illustrate the translation performance on the FLORES-200 benchmark according to varying numbers of in-context exemplars in Figure 2. We can see that the COMET scores increase dramatically with the number of exemplars increasing from to , and the translation performance plateaus afterward except for Mistral-7B-v0.3 which has relatively weak multilingual capability.
5 GemmaX: Boosting Multilingual Translation with Gemma Models
[…]
5.1 Pretraining Data
Monolingual Data
We collect monolingual data from CulturaX Nguyen et al. (2024) and MADLAD-400 Kudugunta et al. (2023). CulturaX contains trillion tokens in languages, which undergo meticulous cleaning and deduplication through a comprehensive pipeline. MADLAD-400 is a multilingual dataset with trillion tokens in languages, on which the authors perform a self-audit to guarantee the data quality.
Parallel Data
We collect all Chinese-centric and English-centric parallel datasets from the OPUS collection111111http://www.opus.nlpl.eu Tiedemann (2012) up to August 2024, which is comprised of multiple corpora, ranging from movie subtitles Tiedemann (2016) to Bible Christodouloupoulos and Steedman (2015) to web crawled datasets El-Kishky et al. (2020); Schwenk et al. (2021). We download all available corpora and concatenate them without curating the datasets or trying to balance the representation of different domains. After collecting all parallel datasets, we adopt the data-cleaning process as follows: 1) We remove duplicate sentence pairs and discard sentence pairs by utilizing some heuristic approaches. 2) Language identification filtering is applied by utilizing the fastText toolkit Joulin et al. (2016, 2017). 3) Semantic similarity filtering is performed based on LaBSE Feng et al. (2022) and MuSR Gao et al. (2023). Specifically, we only keep the sentence pairs that have similarity scores between and . After the cleaning process, we have about 3.4 billion cleaned Chinese-centric and English-centric sentence pairs covering 28 languages. The distribution of our parallel datasets for each language is illustrated in Figure 3.
5.2 Supervised Finetuning Data
Inspired by Xu et al. (2024a) that a small amount of high-quality data could dramatically boost the translation performance, we construct our finetuning dataset mostly from human-annotated records: 1) We extract the sentence pairs covering ten languages from the general translation task in the TowerBlock dataset121212https://huggingface.co/datasets/Unbabel/TowerBlocks-v0.2. 2) For the languages not covered in TowerBlock, we randomly sample English-centric sentence pairs for each language either from the NTREX-128 Federmann et al. (2022) and FLORES-200 dev datasets or the OPUS dataset bidirectionally filtered by wmt23-cometkiwi-da-xxl model with quality scores above . 3) To mitigate the off-target issue and enhance the zero-shot translation performance, inspired by Wu et al. (2024), we also randomly select non-English-centric translation directions and sample sentence pairs for each direction from the NTREX-128 and FLORES-200 dev datasets. In summary, the total number of sentence pairs is around K, where there are about K in the English-centric directions and K in other directions.
5.3 Training Configuration
We train all models with the LLaMA-Factory Zheng et al. (2024) framework for one epoch on and NVIDIA H800 GPUs for the pretraining and finetuning stages respectively.
In the pretraining stage, we utilize an effective batch size of 1.57 million tokens, corresponding to a batch size of per GPU with a gradient accumulation of and a maximum sequence length of . We use the AdamW optimizer with a learning rate of 2e-5. We adopt full-weight pretraining with bf16 precision by employing a cosine learning rate scheduler. In the finetuning stage, we adopt a batch size of per GPU with a gradient accumulation of and a maximum sequence length of . Additionally, we employ an inverse square root learning rate scheduler. Detailed training configurations can be found in Tables 7 and 8. We use greedy decoding to generate translations.
5.4 Exploring the Best Data Recipe for Multilingual Translation with LLMs
Xu et al. (2024a) conduct continual pretraining solely on monolingual data followed by instruction finetuning on a small number of high-quality translation pairs, achieving impressive translation results. They conclude that the reliance on parallel data is diminished in the era of LLMs. Subsequently, Guo et al. (2024) reemphasize the importance of parallel data and propose a three-stage training strategy: 1) continual pretraining with monolingual data; 2) continual pretraining with parallel data; 3) finetuning with source language consistent translation instructions. However, as illustrated in Figure 3, the parallel data adopted in the second stage could be highly imbalanced for massively multilingual settings, which would weaken the model’s understanding and generative capabilities on low-resource languages learned in the first stage. Alves et al. (2024) utilize a multilingual mixture of monolingual (two-thirds) and parallel (one-third) data during continual pretraining, which significantly improves the translation performance of their model. In our experiments, we follow the idea of a multilingual mixture in the continual pretraining stage and raise two questions: 1) whether a monolingual to parallel ratio of 2:1 is the best data recipe for multilingual machine translation? 2) If not, what is the best data recipe?
[…]
We continually pretrain the Gemma2-9B model using five datasets containing varying proportions of monolingual and parallel data, followed by translation instruction finetuning on our high-quality datasets. Based on Table 3, we classify the languages into high-resource (), mid-resource (), and low-resource () languages according to their resource availability. We then report the translation performance during continual pretraining on these three language groups in Figures 4 and 5.
By checking model performance under different combinations of data recipes and language groups, we have the following observations:
The current prevalent training paradigm, which involves continual pretraining on large-scale monolingual data, proves to be suboptimal. While this approach improves the overall performance across various languages, it does so at the cost of translation quality for high-resource languages, particularly in the xxen directions.
Incorporating parallel data into the pretraining stage, regardless of the volume, significantly enhances the model’s overall translation capabilities. Notably, the translation performance for high-resource language directions improves as the quantity of parallel data increases. However, such improvement was not evident for low-resource and mid-resource languages.
We believe that the base model has sufficient understanding and generation capabilities for high-resource languages. It mainly needs parallel data during pretraining to align the internal representations across different languages. In contrast, the model’s generative and comprehension abilities for low-resource and mid-resource languages are not up to par. To improve its performance with these languages, it’s crucial to include more data in the pretraining process to enhance its generative capacities. Unfortunately, due to the nature of low-resource and mid-resource languages, we are unable to obtain enough parallel data. Therefore, we think it’s essential to incorporate monolingual data into the pretraining dataset for low-resource and mid-resource languages.
Model | WMT-24 | FLORES-200 | |||
en xx | en xx | xx en | zh xx | xx zh | |
languages | |||||
TowerInstruct-7B-v0.2 | 84.48 / 75.02 | 38.91 / 88.46 | 42.20 / 88.29 | 23.36 / 85.21 | 26.87 / 85.69 |
TowerInstruct-13B-v0.1 | 86.10 / 76.74 | 40.60 / 88.89 | 43.14 / 88.50 | 25.28 / 86.13 | 29.32 / 86.59 |
GemmaX2-28-2B | 84.50 / 74.42 | 40.15 / 88.56 | 42.03 / 88.27 | 25.71 / 85.85 | 31.91 / 87.04 |
GemmaX2-28-9B | 86.59 / 77.25 | 42.58 / 89.15 | 43.95 / 88.62 | 28.55 / 86.84 | 34.10 / 87.70 |
languages | |||||
Aya-23-8B | 77.72 / 69.55 | 34.95 / 86.81 | 39.56 / 86.57 | 23.68 / 84.28 | 25.33 / 85.15 |
Aya-Expanse-8B | 80.15 / 74.11 | 37.57 / 87.89 | 40.48 / 87.04 | 25.39 / 85.48 | 28.21 / 86.30 |
GemmaX2-28-2B | 80.66 / 73.07 | 39.53 / 88.51 | 42.77 / 88.44 | 25.57 / 85.54 | 31.69 / 86.81 |
GemmaX2-28-9B | 82.95 / 75.82 | 42.27 / 89.21 | 45.23 / 88.92 | 28.68 / 86.67 | 34.43 / 87.59 |
languages | |||||
X-ALMA | 81.67 / 74.73 | 39.31 / 88.69 | 42.09 / 88.54 | - | - |
GemmaX2-28-2B | 79.83 / 72.77 | 39.32 / 88.27 | 42.72 / 88.46 | 25.54 / 85.24 | 31.42 / 86.73 |
GemmaX2-28-9B | 82.05 / 75.31 | 41.98 / 88.97 | 45.27 / 88.97 | 28.64 / 86.36 | 34.26 / 87.56 |
languages | |||||
Aya-101 | 71.67 / 63.88 | 27.70 / 84.80 | 33.84 / 86.21 | 17.36 / 81.56 | 18.70 / 81.72 |
LLaMAX3-Alpaca-8B | 67.29 / 60.21 | 28.11 / 83.12 | 35.81 / 87.05 | 18.17 / 80.78 | 21.08 / 83.38 |
GPT-3.5-turbo | 70.63 / 62.64 | 32.24 / 82.49 | 36.18 / 85.63 | 20.35 / 79.34 | 24.12 / 82.61 |
NLLB-54.5B | - | 37.34 / 87.05 | 43.63 / 88.32 | 25.41 / 84.42 | 20.73 / 80.72 |
GPT-4-turbo | 77.55 / 70.68 | 37.01 / 87.14 | 41.53 / 88.40 | 24.66 / 84.31 | 28.80 / 85.90 |
Google Translate | 77.64 / 73.00 | 41.52 / 88.51 | 45.35 / 88.83 | 29.64 / 85.69 | 34.86 / 87.20 |
GemmaX2-28-2B | 77.20 / 71.99 | 37.00 / 87.54 | 42.16 / 88.32 | 24.27 / 84.44 | 30.59 / 86.41 |
GemmaX2-28-9B | 79.37 / 74.41 | 39.72 / 88.35 | 45.07 / 88.95 | 27.48 / 85.69 | 33.74 / 87.38 |
5.5 Main Result
Based on the experimental results in Section 5.4, we propose a Parallel-First Monolingual-Second (PFMS) data mixing strategy, where we give higher priority to parallel data than monolingual data when preparing the continual pretraining dataset. Specifically, for 2 billion tokens per language, we utilize parallel data as much as possible and supplement it with monolingual data if needed. The detailed information on the number of training tokens across different languages adopted in the PFMS strategy is summarized in Table 9. The translation performance by leveraging the PFMS strategy is also illustrated in Figures 4 and 5. We can see that the PFMS strategy consistently outperforms other data mixing strategies except for high-resource languages to English directions. Such a phenomenon might be due to the fact that the dominant part of the parallel-only dataset is the English-centric corpora with high-resource languages.
We continually pretrain Gemma2-9B with PFMS strategy and learn GemmaX2-28-9B by finetuning the pretrained model with high-quality translation pairs covering languages. Furthermore, to investigate the impact of our PFMS strategy on models of different scales, we also conduct experiments on the Gemma2-2B model. The experimental results are summarized in Table 2. Besides the models discussed in Section 3.2, we include several of the strongest open-source multilingual models as baselines, including:
TowerInstruct-7/13B Alves et al. (2024): 7B and 13B LLaMA2-based models for translation-related tasks supporting languages.
Aya-23-8B / Aya-Expanse-8B Dang et al. (2024): 8B models with highly advanced multilingual capabilities, supporting 23 commonly used languages. Note that we only evaluate their multilingual translation performance on languages overlapping with GemmaX2.
X-ALMA Xu et al. (2024b): a LLaMA2-based multilingual MT model covering languages, which consists of a 13B dense model and multiple language-specific modules, with a total of approximately B parameters. Note that X-ALMA only supports English-centric directions, and we report the translation performance for the overlapping directions.
Aya-101 Üstün et al. (2024): a 13B massively multilingual generative language model that follows instructions in languages.
LLaMAX3-Alpaca-8B Lu et al. (2024): an 8B LLaMA3-based model supporting languages with powerful multilingual capabilities without loss instruction-following capabilities.
Despite having limited parameters and supporting a large number of languages, GemmaX2-28-9B consistently outperforms the current SOTA open-source models, demonstrating the effectiveness of our proposed PFMS data mixing strategy and the powerful translation capabilities of our model. In addition, our model achieves comparable translation performance to GPT-4-turbo and Google Translate, indicating that our model has achieved translation capabilities on par with industry standards. We also observe that GemmaX2-28-2B demonstrates strong multilingual translation capabilities with minimal model parameters, indicating the effectiveness of our PFMS data mixing strategy for models with varying parameter scales.
6 Conclusion
In this paper, we systematically evaluate the multilingual in-context translation of the latest open LLMs at a practical scale. While these models demonstrate strong translation capabilities, they still fall short compared to closed-source models. Furthermore, we explore the best data recipe for multilingual translation with LLMs and propose a Parallel-First Monolingual-Second (PFMS) data mixing strategy. Leveraging the PFMS strategy during continual pretraining on Gemma2-9B, we achieve significant improvements, bringing its translation performance to a level comparable with GPT-4-turbo and Google Translate. To the best of our knowledge, GemmaX2-28-9B is the open model with the highest translation quality. We aim to develop models that support a broader range of languages and possess enhanced translation capabilities as part of our future work.
Limitations
Due to limited computational resources, we only conduct multilingual in-context translation evaluation and explore the optimal data mixing strategy on open LLMs with parameter sizes below ten billion. The translation performance and optimal data recipe for larger models remain unclear.
Appendix A Appendix
ISO Code | Language | Script | Family | Subgrouping | Resource |
ar | Arabic | Arabic | Afro-Asiatic | Semitic | High |
bn | Bengali | Bengali | Indo-European | Indo-Aryan | Mid |
cs | Czech | Latin | Indo-European | Balto-Slavic | High |
de | German | Latin | Indo-European | Germanic | High |
en | English | Latin | Indo-European | Germanic | High |
es | Spanish | Latin | Indo-European | Italic | High |
fa | Persian | Arabic | Indo-European | Iranian | High |
fr | French | Latin | Indo-European | Italic | High |
he | Hebrew | Hebrew | Afro-Asiatic | Semitic | Mid |
hi | Hindi | Devanagari | Indo-European | Indo-Aryan | High |
id | Indonesian | Latin | Austronesian | Malayo-Polynesian | Mid |
it | Italian | Latin | Indo-European | Italic | High |
ja | Japanese | Japanese | Japonic | Japanesic | High |
km | Khmer | Khmer | Austroasiatic | Khmeric | Low |
ko | Korean | Hangul | Koreanic | Korean | High |
lo | Lao | Lao | Tai-Kadai | Kam-Tai | Low |
ms | Malay | Latin | Austronesian | Malayo-Polynesian | Mid |
my | Burmese | Myanmar | Sino-Tibetan | Burmo-Qiangic | Low |
nl | Dutch | Latin | Indo-European | Germanic | High |
pl | Polish | Latin | Indo-European | Balto-Slavic | High |
pt | Portuguese | Latin | Indo-European | Italic | High |
ru | Russian | Cyrillic | Indo-European | Balto-Slavic | High |
th | Thai | Thai | Tai-Kadai | Kam-Tai | Mid |
tl | Tagalog | Latin | Austronesian | Malayo-Polynesian | Mid |
tr | Turkish | Latin | Turkic | Common Turkic | High |
ur | Urdu | Arabic | Indo-European | Indo-Aryan | Mid |
vi | Vietnamese | Latin | Austroasiatic | Vietic | High |
zh | Chinese | Han | Sino-Tibetan | Sinitic | High |
Direction | Mistral-7B | Qwen2-7B | Qwen2.5-7B | Llama3-8B | Llama3.1-8B | Gemma2-9B | GPT3.5-turbo | GPT4-turbo | |
enar | 41.11 / 31.29 | 59.12 / 52.88 | 68.34 / 62.36 | 65.78 / 58.86 | 66.15 / 60.33 | 73.16 / 68.69 | 75.91 / 69.33 | 79.86 / 73.98 | 77.23 / 72.53 |
enbn | 28.37 / 18.93 | 39.30 / 36.79 | 38.51 / 35.22 | 59.02 / 57.20 | 56.73 / 54.98 | 63.46 / 62.29 | 54.49 / 44.36 | 70.77 / 56.86 | 67.36 / 66.82 |
encs | 59.94 / 47.11 | 53.64 / 40.42 | 57.05 / 45.94 | 67.00 / 55.98 | 67.24 / 56.96 | 76.29 / 68.19 | 77.56 / 69.19 | 81.47 / 74.05 | 77.69 / 69.08 |
ende | 85.70 / 62.46 | 84.15 / 60.77 | 86.52 / 65.85 | 87.59 / 68.15 | 87.94 / 68.69 | 90.17 / 73.60 | 91.09 / 75.95 | 92.65 / 77.84 | 93.63 / 79.26 |
enes | 79.67 / 66.42 | 81.22 / 68.50 | 82.19 / 69.88 | 83.19 / 70.74 | 83.07 / 70.57 | 85.25 / 74.24 | 86.66 / 75.96 | 86.79 / 76.47 | 86.32 / 75.20 |
enfa | 28.82 / 18.50 | 47.61 / 42.86 | 44.90 / 39.93 | 65.23 / 61.14 | 65.46 / 61.87 | 75.12 / 72.24 | 68.63 / 63.90 | 76.76 / 72.20 | 73.06 / 69.65 |
enfr | 74.59 / 63.03 | 76.15 / 65.68 | 77.95 / 67.55 | 77.14 / 66.72 | 76.47 / 66.50 | 81.13 / 72.24 | 83.48 / 74.83 | 83.94 / 76.49 | 81.47 / 73.22 |
enhe | 29.06 / 14.47 | 48.81 / 38.16 | 48.24 / 37.52 | 58.44 / 50.15 | 60.01 / 52.30 | 70.53 / 63.78 | 63.21 / 52.77 | 75.65 / 66.95 | 74.62 / 68.29 |
enhi | 32.57 / 26.45 | 37.75 / 34.44 | 37.76 / 35.11 | 57.23 / 55.32 | 58.60 / 56.92 | 62.72 / 62.67 | 58.70 / 47.81 | 68.42 / 54.56 | 77.30 / 71.63 |
enid | 64.89 / 49.71 | 74.99 / 61.66 | 75.48 / 65.49 | 76.96 / 66.28 | 77.13 / 66.55 | 81.35 / 71.79 | 84.71 / 75.28 | 84.43 / 76.72 | 83.11 / 76.00 |
enit | 76.55 / 63.93 | 77.36 / 64.60 | 79.43 / 67.36 | 79.90 / 68.23 | 80.62 / 69.44 | 84.41 / 74.70 | 85.61 / 75.22 | 87.06 / 78.18 | 83.82 / 74.54 |
enja | 59.27 / 61.91 | 71.76 / 74.69 | 73.05 / 76.46 | 72.84 / 74.77 | 72.16 / 74.95 | 80.98 / 81.40 | 81.23 / 78.82 | 83.98 / 83.47 | 86.64 / 86.18 |
enkm | 10.94 / 2.85 | 18.58 / 20.69 | 17.36 / 18.69 | 25.18 / 29.73 | 27.12 / 32.56 | 39.10 / 49.53 | 26.80 / 23.54 | 56.63 / 55.89 | 65.88 / 72.85 |
enko | 59.43 / 57.74 | 67.71 / 65.51 | 71.72 / 70.04 | 73.74 / 69.83 | 72.80 / 70.55 | 76.24 / 71.73 | 79.42 / 74.55 | 84.84 / 80.85 | 80.49 / 78.07 |
enlo | 13.64 / 2.65 | 14.30 / 5.72 | 17.22 / 7.76 | 16.58 / 9.72 | 15.85 / 9.23 | 33.34 / 32.53 | 23.91 / 15.22 | 43.33 / 37.63 | 61.67 / 63.59 |
enms | 60.97 / 46.86 | 70.17 / 57.46 | 70.32 / 60.10 | 72.30 / 62.10 | 72.17 / 62.45 | 76.49 / 66.87 | 82.42 / 72.85 | 82.04 / 74.49 | 76.25 / 67.71 |
enmy | 11.72 / 1.03 | 14.43 / 10.11 | 11.48 / 6.31 | 23.52 / 25.33 | 24.49 / 28.24 | 30.28 / 37.82 | 22.64 / 14.59 | 47.87 / 45.98 | 59.83 / 70.76 |
ennl | 78.58 / 62.17 | 78.46 / 63.04 | 79.07 / 64.96 | 83.81 / 71.15 | 84.39 / 71.31 | 87.27 / 74.96 | 89.65 / 78.53 | 90.16 / 80.58 | 88.51 / 78.86 |
enpl | 66.30 / 52.06 | 63.52 / 48.74 | 65.01 / 53.03 | 72.70 / 60.18 | 73.06 / 60.44 | 81.96 / 71.46 | 81.32 / 71.33 | 85.56 / 76.99 | 82.59 / 74.22 |
enpt | 78.42 / 65.55 | 80.69 / 68.02 | 82.57 / 70.61 | 81.07 / 68.80 | 80.81 / 68.49 | 84.82 / 74.23 | 87.31 / 76.98 | 87.89 / 78.31 | 86.42 / 75.78 |
enru | 71.72 / 61.82 | 75.28 / 67.42 | 76.32 / 68.88 | 75.63 / 67.59 | 76.34 / 67.87 | 80.52 / 74.01 | 80.25 / 72.37 | 83.74 / 76.80 | 81.05 / 74.40 |
enth | 34.48 / 25.08 | 66.46 / 63.25 | 71.25 / 67.34 | 67.95 / 62.92 | 64.96 / 59.28 | 76.44 / 72.16 | 63.23 / 53.22 | 76.65 / 68.99 | 75.48 / 71.67 |
entl | 46.74 / 36.71 | 43.39 / 29.25 | 38.95 / 25.96 | 57.58 / 50.65 | 57.01 / 51.54 | 67.46 / 63.14 | 72.90 / 69.51 | 75.30 / 74.09 | 71.46 / 64.89 |
entr | 40.88 / 31.49 | 47.60 / 42.15 | 51.90 / 49.52 | 60.78 / 57.13 | 63.12 / 60.30 | 72.00 / 69.17 | 75.14 / 70.32 | 77.41 / 75.42 | 77.22 / 75.74 |
enur | 24.00 / 17.08 | 25.92 / 21.18 | 25.30 / 18.12 | 44.41 / 43.99 | 49.00 / 51.48 | 57.71 / 60.42 | 54.85 / 51.51 | 68.77 / 62.92 | 65.71 / 67.00 |
envi | 54.44 / 42.18 | 72.71 / 64.65 | 76.98 / 70.68 | 75.09 / 67.74 | 74.17 / 68.09 | 78.98 / 73.84 | 76.79 / 70.57 | 81.01 / 76.03 | 79.35 / 75.31 |
enzh | 69.74 / 63.57 | 78.18 / 72.73 | 77.92 / 72.35 | 74.85 / 68.23 | 74.32 / 68.26 | 78.34 / 72.52 | 79.09 / 72.65 | 80.98 / 75.69 | 82.04 / 77.64 |
Direction | Mistral-7B | Qwen2-7B | Qwen2.5-7B | Llama3-8B | Llama3.1-8B | Gemma2-9B | NLLB-54B | GPT3.5-turbo | GPT4-turbo | |
aren | 29.99 / 82.55 | 38.92 / 86.46 | 40.60 / 86.94 | 41.42 / 87.01 | 41.94 / 87.01 | 46.04 / 87.94 | 48.26 / 88.13 | 41.83 / 87.77 | 46.39 / 88.44 | 50.59 / 88.73 |
aren | 8.66 / 63.97 | 21.76 / 78.71 | 27.11 / 83.13 | 27.65 / 82.86 | 28.63 / 83.4 | 35.87 / 86.02 | 43.03 / 87.22 | 37.23 / 86.81 | 40.79 / 87.81 | 47.80 / 88.72 |
bnen | 17.28 / 78.07 | 28.01 / 85.6 | 27.96 / 85.75 | 32.74 / 87.25 | 32.42 / 87.15 | 38.52 / 88.61 | 42.22 / 89.31 | 29.17 / 86.37 | 39.19 / 89.32 | 43.52 / 89.62 |
bnen | 3.06 / 44.88 | 9.1 / 66.48 | 8.14 / 64.31 | 21.85 / 80.71 | 21.58 / 80.24 | 23.24 / 82.58 | 36.04 / 87.08 | 19.96 / 78.24 | 29.39 / 86.18 | 37.49 / 86.60 |
csen | 41.61 / 87.91 | 40.78 / 87.52 | 42.18 / 87.87 | 43.39 / 88.23 | 43.87 / 88.39 | 46.53 / 88.96 | 45.31 / 88.67 | 45.09 / 89.02 | 46.44 / 89.28 | 48.68 / 89.34 |
csen | 28.17 / 85.32 | 24.17 / 81.43 | 24.87 / 82.34 | 32.41 / 87.57 | 33.35 / 87.97 | 36.75 / 90.45 | 42.38 / 91.54 | 39.65 / 91.25 | 42.33 / 92.11 | 45.95 / 91.51 |
deen | 46.23 / 88.99 | 45.96 / 88.95 | 45.09 / 88.92 | 47.29 / 88.99 | 47.72 / 89.18 | 50.72 / 89.62 | 49.80 / 89.36 | 48.71 / 89.80 | 50.08 / 89.89 | 52.80 / 90.19 |
deen | 34.35 / 84.6 | 33.12 / 83.78 | 34.51 / 84.81 | 38.54 / 86.06 | 39.39 / 86.2 | 45.01 / 87.88 | 46.62 / 88.10 | 47.10 / 88.52 | 48.67 / 89.09 | 48.27 / 89.29 |
esen | 34.94 / 86.96 | 35.24 / 87.06 | 36.05 / 87.15 | 36.95 / 87.17 | 37.34 / 87.24 | 39.02 / 87.47 | 38.80 / 87.49 | 36.27 / 87.70 | 38.23 / 87.99 | 37.35 / 87.53 |
esen | 28.8 / 84.81 | 29.26 / 85.24 | 30.14 / 85.56 | 30.55 / 85.92 | 30.92 / 85.96 | 33.45 / 86.67 | 33.05 / 86.39 | 34.25 / 87.25 | 34.41 / 87.36 | 35.23 / 87.24 |
faen | 24.54 / 81.42 | 33.47 / 86.21 | 33.80 / 86.12 | 37.92 / 87.55 | 38.19 / 87.58 | 43.07 / 88.65 | 44.19 / 88.70 | 37.35 / 87.83 | 41.58 / 88.82 | 44.68 / 88.90 |
faen | 4.49 / 48.09 | 13.49 / 72.72 | 13.41 / 70.40 | 25.68 / 83.71 | 26.18 / 84.11 | 32.19 / 87.31 | 36.13 / 87.75 | 29.30 / 85.60 | 34.85 / 88.10 | 39.72 / 88.34 |
fren | 47.38 / 89.01 | 47.41 / 89.02 | 47.79 / 89.15 | 49.17 / 89.01 | 49.57 / 89.19 | 51.93 / 89.63 | 51.54 / 89.44 | 49.93 / 89.70 | 51.41 / 89.88 | 53.44 / 89.89 |
fren | 45.66 / 85.95 | 46.32 / 86.59 | 47.23 / 86.80 | 49.43 / 87.13 | 50.12 / 87.2 | 55.12 / 88.36 | 56.16 / 88.15 | 56.53 / 89.09 | 56.59 / 89.13 | 58.98 / 89.23 |
heen | 24.07 / 76.46 | 40.28 / 86.57 | 41.28 / 86.70 | 44.98 / 87.81 | 45.02 / 87.84 | 49.33 / 88.80 | 49.02 / 88.62 | 42.60 / 87.57 | 47.73 / 88.99 | 51.80 / 89.19 |
heen | 3.99 / 46.57 | 17.12 / 74.23 | 15.95 / 71.86 | 30.32 / 83.23 | 30.67 / 83.55 | 34.09 / 85.84 | 46.82 / 88.71 | 33.29 / 83.78 | 41.89 / 88.58 | 48.80 / 88.88 |
hien | 23.34 / 82.32 | 33.57 / 87.3 | 33.87 / 87.12 | 39.15 / 88.97 | 39.25 / 88.97 | 45.56 / 89.79 | 47.26 / 90.27 | 37.57 / 88.69 | 44.75 / 90.07 | 51.52 / 90.99 |
hien | 6.32 / 49.27 | 11.79 / 61.07 | 12.18 / 61.34 | 25.77 / 75 | 27.56 / 75.92 | 35.33 / 79.07 | 40.55 / 81.17 | 27.71 / 76.49 | 36.37 / 81.01 | 38.94 / 83.13 |
iden | 41.26 / 87.9 | 44.26 / 88.94 | 45.78 / 89.16 | 46.41 / 89.16 | 46.72 / 89.29 | 50.87 / 90.00 | 49.88 / 89.37 | 47.66 / 89.83 | 49.72 / 90.22 | 53.25 / 90.29 |
iden | 24.9 / 82.61 | 34.75 / 88.5 | 37.40 / 89.19 | 41.02 / 90.06 | 41.4 / 90.15 | 48.53 / 91.55 | 49.18 / 91.18 | 47.13 / 91.80 | 49.41 / 92.29 | 52.76 / 92.61 |
iten | 38.18 / 87.76 | 38.32 / 87.82 | 38.89 / 87.90 | 39.96 / 87.93 | 40.1 / 88.1 | 41.9 / 88.52 | 41.56 / 88.13 | 39.35 / 88.45 | 40.58 / 88.67 | 41.31 / 88.54 |
iten | 30.19 / 86.13 | 28.35 / 85.71 | 29.08 / 86.33 | 32.98 / 87.39 | 33.63 / 87.53 | 36.39 / 88.56 | 38.27 / 88.58 | 37.80 / 89.13 | 38.77 / 89.34 | 39.86 / 89.38 |
jaen | 26.03 / 85.79 | 30.11 / 87.43 | 31.01 / 87.64 | 30.89 / 87.59 | 30.81 / 87.5 | 34.51 / 88.24 | 34.54 / 87.92 | 30.53 / 88.24 | 33.61 / 88.68 | 38.14 / 89.36 |
jaen | 13.25 / 82.79 | 21.12 / 88.66 | 24.30 / 89.34 | 24.45 / 89.04 | 24.54 / 89.01 | 30.4 / 91.03 | 20.07 / 89.08 | 30.50 / 90.91 | 32.46 / 91.67 | 37.08 / 92.75 |
kmen | 7.82 / 60.54 | 16.21 / 77.29 | 14.79 / 74.12 | 23.86 / 82.01 | 23.98 / 82.56 | 30.3 / 84.77 | 38.64 / 87.02 | 13.09 / 73.18 | 32.98 / 87.01 | 35.10 / 85.82 |
kmen | 0.19 / 30.09 | 1.9 / 45.18 | 1.59 / 41.26 | 4.82 / 56.06 | 5.44 / 57.08 | 7.15 / 66.08 | 22.97 / 79.86 | 4.12 / 49.19 | 17.93 / 79.92 | 27.36 / 83.37 |
koen | 26.42 / 85.78 | 30.3 / 87.18 | 31.78 / 87.59 | 31.27 / 87.24 | 31.75 / 87.34 | 35.96 / 88.47 | 35.42 / 88.03 | 32.12 / 88.12 | 36.07 / 89.06 | 39.31 / 89.32 |
koen | 10.1 / 79.37 | 15.49 / 85.14 | 17.71 / 86.23 | 19.31 / 86.46 | 20.21 / 86.98 | 23.24 / 88.1 | 26.72 / 89.47 | 24.33 / 88.83 | 28.56 / 90.29 | 31.31 / 90.24 |
loen | 4.92 / 53.28 | 9.64 / 64.26 | 11.03 / 64.81 | 14.57 / 71.85 | 15.17 / 72.33 | 27.13 / 80.54 | 42.40 / 87.67 | 8.29 / 64.60 | 23.41 / 79.64 | 44.05 / 88.53 |
loen | 0.1 / 31.6 | 0.86 / 35.42 | 0.92 / 33.22 | 1.49 / 36.92 | 1.47 / 37.9 | 5.9 / 55.2 | 29.59 / 84.10 | 2.76 / 42.76 | 11.24 / 63.45 | 29.59 / 83.19 |
msen | 40.32 / 87.03 | 42.12 / 87.84 | 42.88 / 87.87 | 46.05 / 88.71 | 46.37 / 88.66 | 51.42 / 89.57 | 51.20 / 89.21 | 47.75 / 89.24 | 50.22 / 89.87 | 53.29 / 89.76 |
msen | 20.16 / 79.33 | 25.95 / 84.98 | 25.08 / 85.03 | 35.99 / 87.29 | 36.34 / 87.29 | 41.95 / 89.15 | 45.54 / 89.01 | 37.36 / 89.39 | 43.86 / 90.14 | 47.41 / 89.91 |
myen | 2.22 / 54.38 | 6.12 / 69.18 | 4.63 / 64.70 | 17.69 / 81.08 | 18.44 / 81.62 | 24.75 / 84.00 | 34.73 / 87.50 | 2.27 / 58.68 | 25.87 / 85.72 | 34.97 / 87.23 |
myen | 0.09 / 31.99 | 1.39 / 42.81 | 0.74 / 36.87 | 6.63 / 64 | 7.13 / 65.03 | 6.05 / 69.63 | 17.74 / 84.58 | 1.39 / 48.20 | 12.85 / 79.26 | 24.45 / 87.50 |
nlen | 35.02 / 86.93 | 35.43 / 87.06 | 35.24 / 87.15 | 35.87 / 86.67 | 36.32 / 87.1 | 38.14 / 87.65 | 38.94 / 87.62 | 37.82 / 88.05 | 38.86 / 88.20 | 39.68 / 87.97 |
nlen | 27.82 / 84.88 | 26.23 / 84.2 | 26.11 / 84.45 | 30.49 / 86.64 | 31 / 86.63 | 33.48 / 87.67 | 35.61 / 87.74 | 36.79 / 88.79 | 37.48 / 88.90 | 37.58 / 88.72 |
plen | 33.71 / 85.8 | 33.58 / 85.71 | 33.86 / 85.75 | 34.75 / 85.85 | 34.82 / 85.9 | 37.1 / 86.76 | 36.68 / 86.35 | 35.45 / 86.66 | 36.38 / 86.97 | 38.05 / 86.93 |
plen | 23.58 / 84.69 | 19.87 / 81.34 | 20.61 / 82.11 | 26.18 / 86.96 | 26.14 / 87.1 | 29.78 / 89.07 | 32.54 / 89.38 | 31.75 / 89.67 | 33.43 / 90.36 | 36.69 / 90.28 |
pten | 50.5 / 89.19 | 51.66 / 89.29 | 52.42 / 89.43 | 52.86 / 89.34 | 53.35 / 89.46 | 56.49 / 90.06 | 55.19 / 89.57 | 54.20 / 90.04 | 55.47 / 90.15 | 57.44 / 90.30 |
pten | 45.43 / 87.36 | 45.59 / 88.17 | 47.61 / 88.57 | 49.34 / 88.67 | 49.46 / 88.71 | 54.04 / 89.78 | 52.87 / 89.26 | 55.85 / 90.42 | 56.95 / 90.57 | 56.98 / 90.58 |
ruen | 38.65 / 86.25 | 38.25 / 86.33 | 39.07 / 86.63 | 39.68 / 86.37 | 39.9 / 86.48 | 42.51 / 87.06 | 42.21 / 86.98 | 39.18 / 87.04 | 41.04 / 87.27 | 43.95 / 87.52 |
ruen | 31.17 / 86.29 | 31.25 / 86.93 | 32.42 / 87.61 | 34.36 / 87.82 | 34.88 / 87.77 | 38.61 / 89.46 | 41.03 / 89.45 | 37.86 / 89.51 | 40.71 / 90.36 | 44.35 / 90.60 |
then | 20.03 / 80.62 | 31.91 / 87.25 | 33.40 / 87.76 | 33.91 / 87.68 | 34.36 / 87.71 | 37.92 / 88.59 | 36.87 / 87.88 | 27.97 / 86.70 | 35.76 / 88.90 | 36.96 / 88.07 |
then | 6.01 / 55.55 | 27.29 / 83.73 | 31.36 / 85.11 | 31.13 / 84.93 | 30.69 / 83.82 | 37.71 / 87.72 | 35.08 / 85.71 | 31.57 / 84.16 | 40.68 / 88.65 | 46.31 / 88.77 |
tlen | 36.99 / 82.95 | 40.03 / 84.37 | 38.55 / 83.14 | 45.97 / 86.24 | 46.36 / 86.45 | 52.44 / 88.01 | 54.58 / 88.38 | 48.54 / 87.93 | 53.96 / 88.87 | 57.18 / 88.78 |
tlen | 11.21 / 66.19 | 11.08 / 67.32 | 8.10 / 61.47 | 25.45 / 79.77 | 26.19 / 79.67 | 33.09 / 83.87 | 38.34 / 84.79 | 33.79 / 84.99 | 39.84 / 86.13 | 39.76 / 84.79 |
tren | 28.22 / 84.38 | 33.75 / 86.91 | 34.72 / 87.38 | 38.42 / 88.64 | 38.73 / 88.6 | 43.68 / 89.68 | 45.82 / 89.84 | 42.29 / 89.86 | 44.80 / 90.23 | 48.79 / 90.48 |
tren | 8.35 / 65.91 | 13.06 / 77.01 | 18.33 / 80.25 | 24.86 / 84.88 | 25.52 / 85.26 | 34.7 / 88.63 | 41.52 / 89.84 | 37.20 / 89.53 | 40.32 / 90.58 | 45.38 / 91.25 |
uren | 14.07 / 73.73 | 25.92 / 83.11 | 24.44 / 82.00 | 31.89 / 85.57 | 32.82 / 85.98 | 38.73 / 87.45 | 43.08 / 88.30 | 30.82 / 85.71 | 39.96 / 88.54 | 43.26 / 88.39 |
uren | 1.83 / 39.69 | 3.59 / 53.02 | 2.30 / 46.69 | 13.05 / 71.02 | 16.64 / 74.04 | 20.58 / 78.4 | 30.52 / 81.75 | 20.20 / 76.07 | 28.71 / 82.92 | 32.46 / 83.05 |
vien | 31.87 / 84.55 | 38.52 / 87.11 | 39.39 / 87.43 | 39.79 / 87.14 | 40.49 / 87.37 | 43.55 / 88.04 | 43.75 / 87.79 | 38.48 / 87.62 | 41.87 / 88.34 | 45.40 / 88.40 |
vien | 18.83 / 74.06 | 34.62 / 86.56 | 37.17 / 87.84 | 38.24 / 87.46 | 38.37 / 87.66 | 42.31 / 89.01 | 43.30 / 88.19 | 39.36 / 88.50 | 43.07 / 89.63 | 47.08 / 89.90 |
zhen | 29.33 / 85.88 | 34.2 / 87.22 | 34.80 / 87.23 | 32.66 / 86.77 | 32.66 / 86.86 | 35.96 / 87.52 | 36.13 / 87.16 | 32.60 / 87.48 | 34.97 / 87.72 | 39.82 / 88.46 |
zhen | 23.28 / 83.37 | 37.11 / 88.43 | 37.06 / 88.36 | 32.2 / 86.99 | 32.5 / 87.04 | 37.28 / 88.42 | 26.61 / 82.25 | 35.72 / 88.42 | 37.84 / 88.96 | 43.36 / 89.94 |
Direction | Mistral-7B | Qwen2-7B | Qwen2.5-7B | Llama3-8B | Llama3.1-8B | Gemma2-9B | NLLB-54B | GPT3.5-turbo | GPT4-turbo | |
arzh | 6.78 / 67.75 | 26.64 / 84.24 | 27.49 / 84.64 | 22.01 / 82.54 | 22.84 / 82.66 | 28.54 / 84.87 | 21.25 / 80.25 | 25.05 / 83.20 | 29.73 / 85.48 | 36.21 / 86.50 |
arzh | 2.91 / 56.15 | 14.7 / 77.21 | 16.53 / 80.10 | 14.09 / 78.08 | 15.96 / 79.07 | 22.66 / 82.16 | 27.53 / 84.08 | 22.66 / 83.65 | 26.24 / 84.62 | 32.32 / 85.30 |
bnzh | 5.16 / 65.41 | 21.86 / 82.77 | 21.24 / 82.69 | 20.77 / 82.3 | 21.35 / 82.68 | 25.39 / 84.94 | 20.01 / 81.33 | 19.40 / 81.58 | 27.38 / 86.37 | 33.55 / 87.07 |
bnzh | 0.74 / 37.07 | 5.54 / 61.44 | 4.85 / 58.94 | 12.81 / 74.3 | 12.71 / 73.98 | 13.27 / 76.97 | 23.67 / 81.96 | 12.81 / 72.07 | 20.64 / 82.36 | 27.64 / 82.38 |
cszh | 17.4 / 80.41 | 28.93 / 85.82 | 28.40 / 85.64 | 25.2 / 84.51 | 25.62 / 84.74 | 30.42 / 86.25 | 19.60 / 79.15 | 29.11 / 86.06 | 31.30 / 86.62 | 36.19 / 87.51 |
cszh | 11.15 / 75.53 | 14.45 / 79.42 | 14.07 / 80.08 | 18.13 / 85.07 | 19.24 / 86.15 | 22.17 / 87.39 | 25.63 / 88.81 | 24.22 / 89.02 | 26.90 / 90.42 | 30.65 / 89.81 |
dezh | 19.32 / 81.31 | 30.96 / 86.8 | 30.68 / 86.81 | 27.33 / 85.5 | 27.42 / 85.51 | 31.82 / 86.93 | 21.48 / 80.87 | 30.38 / 86.82 | 31.65 / 87.07 | 37.03 / 87.97 |
dezh | 16.15 / 77.83 | 20.08 / 80.72 | 21.06 / 81.75 | 21.2 / 82.14 | 22.08 / 82.5 | 24.65 / 83.03 | 27.15 / 84.27 | 27.06 / 84.84 | 29.73 / 85.93 | 32.97 / 86.27 |
enzh | 23.28 / 83.37 | 37.11 / 88.43 | 37.06 / 88.36 | 32.2 / 86.99 | 32.5 / 87.04 | 37.28 / 88.42 | 26.61 / 82.25 | 35.72 / 88.42 | 37.97 / 88.51 | 43.36 / 89.94 |
enzh | 29.33 / 85.88 | 34.2 / 87.22 | 34.80 / 87.23 | 32.66 / 86.77 | 32.66 / 86.86 | 35.96 / 87.52 | 36.13 / 87.16 | 32.60 / 87.48 | 35.23 / 87.78 | 39.82 / 88.46 |
eszh | 16.91 / 81.54 | 27.93 / 86.63 | 28.40 / 86.76 | 24.26 / 85.33 | 24.77 / 85.61 | 28.02 / 86.81 | 19.03 / 79.85 | 27.60 / 86.63 | 28.42 / 86.74 | 32.74 / 87.24 |
eszh | 15.06 / 80.17 | 20.56 / 83.72 | 20.93 / 83.94 | 20.09 / 83.44 | 20.65 / 83.64 | 23.26 / 84.64 | 23.69 / 84.46 | 23.40 / 85.35 | 24.04 / 85.35 | 25.59 / 85.47 |
fazh | 9.14 / 71.62 | 24.98 / 84.44 | 24.60 / 84.16 | 23.02 / 83.81 | 23.15 / 84.16 | 27.38 / 85.66 | 21.89 / 81.66 | 24.32 / 84.47 | 28.84 / 86.17 | 34.43 / 87.16 |
fazh | 1.67 / 43.92 | 9.43 / 71 | 8.21 / 68.32 | 14.84 / 79.92 | 16.61 / 81.01 | 21.16 / 83.97 | 23.10 / 84.43 | 18.45 / 82.34 | 23.34 / 85.83 | 27.68 / 85.42 |
frzh | 19.01 / 81.3 | 31.89 / 87 | 30.97 / 86.91 | 26.6 / 85.4 | 27.07 / 85.6 | 31.56 / 86.88 | 21.19 / 80.44 | 30.31 / 86.79 | 32.32 / 87.44 | 37.71 / 87.90 |
frzh | 22.32 / 79.96 | 28.14 / 83.33 | 28.62 / 83.24 | 26.7 / 82.68 | 27.68 / 83.04 | 28.42 / 81.98 | 34.16 / 84.28 | 32.94 / 85.22 | 34.58 / 85.73 | 38.24 / 85.89 |
hezh | 7.65 / 66.82 | 26.08 / 83.58 | 25.99 / 83.70 | 23.36 / 82.85 | 24.38 / 83.12 | 29.31 / 85.17 | 19.77 / 78.50 | 24.92 / 83.29 | 30.47 / 85.66 | 36.89 / 86.80 |
hezh | 1.09 / 41.27 | 8.92 / 71.81 | 7.65 / 69.60 | 16.25 / 78.76 | 16.4 / 79.35 | 18.61 / 81.91 | 27.04 / 84.87 | 17.26 / 79.40 | 24.25 / 85.04 | 31.91 / 85.77 |
hizh | 7.18 / 69.31 | 24.34 / 83.71 | 23.66 / 83.41 | 23.35 / 83.75 | 23.58 / 84.09 | 27.86 / 85.47 | 21.12 / 81.60 | 23.27 / 83.75 | 29.00 / 86.48 | 35.64 / 87.51 |
hizh | 1.93 / 40.23 | 6.12 / 52.79 | 7.02 / 54.20 | 14.35 / 67.57 | 14.97 / 67.84 | 17.84 / 70.81 | 24.20 / 73.81 | 16.06 / 69.22 | 22.60 / 74.71 | 28.80 / 75.99 |
idzh | 16.48 / 79.47 | 30.22 / 86.13 | 30.57 / 86.24 | 26.25 / 84.81 | 26.79 / 85.12 | 31.15 / 86.49 | 21.38 / 79.92 | 28.83 / 86.10 | 31.36 / 86.56 | 37.24 / 87.66 |
idzh | 7.72 / 75.08 | 20.73 / 85.59 | 22.22 / 86.64 | 21.6 / 86.24 | 22.62 / 86.77 | 26.44 / 87.2 | 29.55 / 87.99 | 25.70 / 88.21 | 28.86 / 89.22 | 33.18 / 89.40 |
itzh | 17.95 / 81.48 | 29.17 / 87.01 | 29.20 / 86.70 | 24.87 / 85.33 | 25.51 / 85.47 | 29.14 / 86.83 | 19.33 / 79.68 | 28.32 / 86.74 | 29.77 / 87.17 | 34.59 / 87.65 |
itzh | 14.88 / 80.28 | 19.26 / 84.04 | 19.42 / 84.48 | 20.1 / 84.58 | 21.04 / 84.9 | 23.9 / 85.94 | 25.77 / 85.73 | 25.47 / 86.99 | 27.39 / 87.63 | 28.98 / 87.31 |
jazh | 15.36 / 81.36 | 25.74 / 87.51 | 26.62 / 87.54 | 21.79 / 85.61 | 22.32 / 85.9 | 26.06 / 87.4 | 19.64 / 81.81 | 24.94 / 87.42 | 26.09 / 87.16 | 33.54 / 88.71 |
jazh | 8.58 / 81.35 | 16.65 / 87.72 | 18.59 / 88.68 | 17.42 / 87.32 | 18.15 / 87.69 | 21.51 / 89.08 | 16.78 / 87.83 | 22.44 / 89.74 | 24.42 / 90.76 | 28.21 / 90.97 |
kmzh | 1.26 / 49.34 | 13.74 / 76.05 | 10.52 / 71.66 | 14.78 / 78.29 | 16.12 / 78.85 | 20.82 / 82.03 | 21.33 / 83.00 | 9.16 / 69.56 | 23.85 / 84.61 | 29.39 / 85.41 |
kmzh | 0.14 / 28.15 | 1.47 / 42.66 | 1.38 / 39.96 | 3.25 / 51.52 | 3.81 / 52.31 | 4.51 / 60.19 | 17.55 / 79.83 | 2.83 / 46.69 | 14.22 / 77.03 | 22.26 / 80.68 |
kozh | 14.79 / 79.61 | 26.7 / 86.41 | 27.12 / 86.50 | 23.4 / 84.71 | 24.23 / 85.17 | 26.74 / 86.19 | 21.31 / 83.01 | 25.00 / 85.82 | 28.56 / 86.89 | 34.03 / 87.87 |
kozh | 4.39 / 73.84 | 11.77 / 83.31 | 13.79 / 85.00 | 14.25 / 84.34 | 15.65 / 85.14 | 16.98 / 85.8 | 20.39 / 86.86 | 17.77 / 86.67 | 21.65 / 88.40 | 24.80 / 88.12 |
lozh | 0.94 / 46.14 | 6.6 / 62.99 | 8.62 / 63.95 | 8.02 / 68.08 | 9.25 / 69.03 | 16.84 / 77.22 | 23.44 / 83.71 | 5.57 / 61.33 | 16.28 / 76.24 | 33.45 / 86.67 |
lozh | 0.09 / 30.12 | 0.51 / 34.27 | 0.58 / 32.13 | 0.74 / 34.22 | 0.89 / 34.62 | 2.67 / 47.97 | 24.30 / 82.08 | 1.50 / 41.01 | 7.66 / 59.23 | 22.51 / 80.22 |
mszh | 15.32 / 78.42 | 27.99 / 84.8 | 28.33 / 84.74 | 24.98 / 83.76 | 26.02 / 84.16 | 30.16 / 85.66 | 23.16 / 80.89 | 28.11 / 85.11 | 30.49 / 85.89 | 36.75 / 87.02 |
mszh | 6.28 / 72 | 14.14 / 82.04 | 14.67 / 82.71 | 18.48 / 83.34 | 19.16 / 83.82 | 21.86 / 84.93 | 26.33 / 85.77 | 18.77 / 85.32 | 23.98 / 86.76 | 30.23 / 86.50 |
myzh | 0.21 / 43.3 | 5.8 / 69.34 | 3.99 / 63.66 | 12.51 / 77.49 | 13.76 / 78.56 | 16.98 / 80.35 | 16.77 / 80.42 | 1.27 / 55.77 | 18.91 / 82.18 | 27.72 / 85.02 |
myzh | 0.02 / 29.98 | 1.2 / 40.85 | 0.68 / 34.90 | 4.72 / 56.48 | 5.16 / 58.1 | 3.88 / 62.3 | 15.79 / 82.04 | 1.20 / 45.56 | 10.39 / 75.62 | 20.33 / 84.40 |
nlzh | 15.9 / 80.08 | 27.38 / 85.72 | 27.21 / 85.61 | 23.79 / 84.21 | 24.22 / 84.54 | 27.29 / 85.8 | 16.61 / 77.64 | 26.87 / 85.91 | 28.46 / 86.32 | 32.10 / 86.66 |
nlzh | 13.09 / 77.7 | 17.57 / 81.87 | 17.77 / 82.17 | 19.2 / 82.99 | 19.78 / 83.7 | 21.48 / 84.03 | 24.74 / 85.31 | 23.69 / 85.53 | 25.89 / 86.64 | 28.00 / 86.50 |
plzh | 16.04 / 79.7 | 26.73 / 85.3 | 26.58 / 85.14 | 22.87 / 83.61 | 23.81 / 84.04 | 27.42 / 85.44 | 17.37 / 77.75 | 25.96 / 85.12 | 28.00 / 85.76 | 32.27 / 86.52 |
plzh | 13.15 / 78.46 | 15.14 / 81.76 | 14.70 / 81.59 | 17.6 / 85.21 | 18.1 / 85.73 | 21 / 86.86 | 23.19 / 87.82 | 22.94 / 88.56 | 25.06 / 89.37 | 27.95 / 89.29 |
ptzh | 18.81 / 81.45 | 31.29 / 87.27 | 31.02 / 87.15 | 26.51 / 85.63 | 27.47 / 85.94 | 31.31 / 87.25 | 20.54 / 80.02 | 30.22 / 86.57 | 32.78 / 87.46 | 38.01 / 88.43 |
ptzh | 19.15 / 81.99 | 25.5 / 85.6 | 27.19 / 85.78 | 25.14 / 85.13 | 25.98 / 85.54 | 29.62 / 86.11 | 31.17 / 86.39 | 29.93 / 87.28 | 32.44 / 87.66 | 34.84 / 87.67 |
ruzh | 15.22 / 78.65 | 29.07 / 85.74 | 29.43 / 85.88 | 24.57 / 83.81 | 25.1 / 84.1 | 29.61 / 85.43 | 21.22 / 80.67 | 27.55 / 85.34 | 29.77 / 85.75 | 35.13 / 86.64 |
ruzh | 13.36 / 75.88 | 20.64 / 85.83 | 22.11 / 86.42 | 20.91 / 85.79 | 21.69 / 86.52 | 24.22 / 86.68 | 28.21 / 87.99 | 24.58 / 88.01 | 27.23 / 88.47 | 31.52 / 88.99 |
thzh | 4.48 / 65.73 | 26.43 / 86.56 | 27.00 / 86.69 | 22.56 / 84.69 | 23.2 / 85.15 | 27.49 / 86.46 | 20.07 / 82.01 | 20.47 / 83.73 | 28.23 / 87.27 | 31.78 / 87.03 |
thzh | 2.15 / 55.72 | 21.29 / 81.02 | 24.81 / 83.01 | 24.55 / 82.23 | 22.19 / 80.98 | 29.39 / 85 | 30.20 / 84.27 | 24.51 / 80.97 | 33.01 / 86.53 | 38.26 / 86.75 |
tlzh | 13.26 / 73.48 | 25.35 / 81.22 | 23.90 / 80.34 | 22.66 / 80.99 | 23.82 / 81.47 | 29.7 / 84.14 | 21.61 / 79.12 | 25.58 / 82.62 | 31.03 / 84.90 | 37.18 / 86.02 |
tlzh | 3.51 / 60.23 | 4.86 / 62.35 | 3.51 / 57.10 | 10.59 / 73.37 | 12.24 / 74.61 | 19.13 / 80.13 | 22.81 / 80.87 | 19.59 / 81.25 | 23.29 / 82.36 | 25.30 / 81.40 |
trzh | 13.63 / 76.11 | 26.32 / 83.84 | 26.51 / 84.42 | 23.87 / 83.59 | 24.54 / 83.95 | 28.63 / 85.39 | 20.27 / 79.47 | 27.21 / 85.15 | 30.30 / 86.38 | 36.06 / 87.42 |
trzh | 3.58 / 59.25 | 10.44 / 74.55 | 10.88 / 75.85 | 13.37 / 79.38 | 16.72 / 80.6 | 20.64 / 82.8 | 24.82 / 84.47 | 21.99 / 84.55 | 23.67 / 85.79 | 30.19 / 86.85 |
urzh | 2.77 / 58.04 | 19.76 / 80 | 18.69 / 79.35 | 20.07 / 81.06 | 21.24 / 81.99 | 24.59 / 83.9 | 21.79 / 82.21 | 20.06 / 81.23 | 27.07 / 85.45 | 33.02 / 86.07 |
urzh | 0.59 / 34.17 | 2.14 / 48.52 | 1.17 / 40.56 | 6.72 / 63.72 | 8.66 / 68.09 | 11.31 / 73.18 | 20.39 / 78.26 | 11.72 / 70.27 | 19.34 / 79.05 | 23.44 / 78.96 |
vizh | 14.77 / 79.62 | 28.11 / 86.63 | 29.17 / 86.72 | 24.38 / 85.22 | 25.15 / 85.62 | 28.37 / 86.69 | 21.79 / 82.34 | 26.03 / 86.01 | 29.70 / 86.75 | 35.14 / 87.95 |
vizh | 7.88 / 70.03 | 24.52 / 85.44 | 26.99 / 87.00 | 24.78 / 85.28 | 25.94 / 85.99 | 28.15 / 86.12 | 31.75 / 87.63 | 27.31 / 87.10 | 29.76 / 88.12 | 34.69 / 88.74 |
Training batch size on each device | 4 |
Number of GPUs | 32 |
Gradient Accumulation Steps | 6 |
Maximum Sequence Length | 2048 |
Number of Epochs | 1 |
Learning Rate | 2e-5 |
LR Scheduler | cosine |
Finetuning Type | full |
bf16 | true |
Template | empty |
Maximum Gradient Norm | 1.0 |
Warmup | 0.01 |
Weight Decay | 0.01 |
Optimizer | AdamW |
DeepSpeed | ZeRO2 |
Training batch size on each device | 4 |
Number of GPUs | 8 |
Gradient Accumulation Steps | 8 |
Maximum Sequence Length | 2048 |
Number of Epochs | 1 |
Learning Rate | 2e-5 |
LR Scheduler | inverse sqrt |
Finetuning Type | full |
bf16 | true |
Template | empty |
Maximum Gradient Norm | 1.0 |
Warmup | 0.01 |
Weight Decay | 0.01 |
Optimizer | AdamW |
DeepSpeed | ZeRO2 |
Languages | # Monolingual tokens | # Parallel tokens (English-centric) | # Parallel tokens (Chinese-centric) |
ar | 0 | 1000000000 | 1000000000 |
bn | 504825391 | 1452758417 | 42416192 |
cs | 0 | 1712873789 | 287126211 |
de | 0 | 1325095947 | 674904053 |
en | 0 | - | 2000000000 |
es | 0 | 1000000000 | 1000000000 |
fa | 1114273750 | 823160488 | 62565762 |
fr | 0 | 1000000000 | 1000000000 |
he | 332751438 | 1526487076 | 140761486 |
hi | 0 | 1932427027 | 67572973 |
id | 0 | 1800539057 | 199460943 |
it | 0 | 1506228911 | 493771089 |
ja | 0 | 1679229006 | 320770994 |
km | 1883737183 | 113850449 | 2412368 |
ko | 815056806 | 964279235 | 220663959 |
lo | 1154721722 | 68003946 | 262157 |
ms | 1435182152 | 544982172 | 19835676 |
my | 1832116262 | 160466198 | 7417540 |
nl | 0 | 1676682297 | 323317703 |
pl | 0 | 1720056384 | 279943616 |
pt | 0 | 1385756006 | 614243994 |
ru | 0 | 1000000000 | 1000000000 |
th | 1461701620 | 511027417 | 27270963 |
tl | 1090609718 | 898064981 | 11325301 |
tr | 0 | 1765499728 | 234500272 |
ur | 1393220838 | 600203213 | 6575949 |
vi | 0 | 1714651057 | 285348943 |
zh | 0 | 2000000000 | - |
Direction | TowerInstruct-7B | TowerInstruct-13B | X-ALMA | Aya-101 | LLaMAX3-Alpaca-8B | GemmaX2-28-2B | GemmaX2-28-9B |
enar | - | - | 78.72 / 73.56 | 67.77 / 62.85 | 65.41 / 58.95 | 75.26 / 70.21 | 78.16 / 73.06 |
enbn | - | - | - | 64.62 / 60.78 | 58.3 / 56.21 | 70.93 / 70.36 | 70.53 / 70.86 |
encs | - | - | 80.1 / 72.76 | 69.62 / 61.35 | 66.97 / 57.09 | 77.75 / 69.11 | 81.48 / 73.48 |
ende | 91.05 / 74.12 | 91.99 / 76.28 | 92.64 / 76.6 | 86.69 / 65.64 | 86.34 / 65.96 | 90.84 / 73.67 | 92.42 / 76.63 |
enes | 86.46 / 75.03 | 87.44 / 75.97 | 86.93 / 76.16 | 80.22 / 67.42 | 81.36 / 67.9 | 86.45 / 74.22 | 88.05 / 76.96 |
enfa | - | - | 78.59 / 74.49 | 68.46 / 65.64 | 65.79 / 62.31 | 76.22 / 72.41 | 79.84 / 75.87 |
enfr | 83.01 / 73.74 | 84.02 / 75.33 | 83.05 / 75.1 | 74.04 / 62.09 | 74.28 / 63.05 | 82.07 / 72.28 | 84.46 / 75.66 |
enhe | - | - | 79.68 / 72.82 | 66.34 / 58.6 | 63.9 / 56.43 | 77.87 / 71.03 | 81.12 / 74.65 |
enhi | - | - | 70.12 / 64.37 | 61.31 / 54.32 | 56.35 / 55.55 | 65.08 / 64.38 | 66.18 / 66.03 |
enid | - | - | 84.53 / 76.62 | 77.79 / 67.8 | 76.17 / 63.08 | 82.22 / 73.63 | 83.58 / 74.95 |
enit | 85.87 / 75.57 | 86.91 / 76.71 | 85.53 / 76.12 | 77.5 / 63.45 | 77.78 / 66.23 | 85.69 / 74.16 | 87.1 / 76.89 |
enja | - | - | 82.54 / 81.23 | 73.32 / 73.39 | 71.6 / 74.13 | 79.59 / 79.01 | 81.81 / 80.82 |
enkm | - | - | - | 67.42 / 72.55 | 45.71 / 52.32 | 65.64 / 71.71 | 68.21 / 74.38 |
enko | 80.87 / 77.51 | 83.83 / 80.01 | 82.03 / 77.65 | 68.96 / 65.74 | 74.22 / 69.81 | 80.52 / 77.34 | 84.73 / 80.8 |
enlo | - | - | - | 69.68 / 67.24 | 43.31 / 32.9 | 62.3 / 64.41 | 64.35 / 66.5 |
enms | - | - | 81.13 / 73.29 | 76.69 / 61.22 | 72.73 / 61.18 | 78.96 / 69.98 | 78.29 / 69.64 |
enmy | - | - | - | 58.72 / 67.7 | 37.44 / 41.92 | 61.72 / 71.46 | 67.37 / 75.74 |
ennl | 89.81 / 78.88 | 90.75 / 79.7 | 90.2 / 79.38 | 83.16 / 68.3 | 81.09 / 68.13 | 88.89 / 77 | 91.07 / 80.18 |
enpl | - | - | 83.93 / 74.21 | 74.44 / 62.37 | 70.87 / 59.21 | 81.29 / 69.89 | 84.64 / 74.18 |
enpt | 86.16 / 76.24 | 87.98 / 77.31 | 86.72 / 77.18 | 81.39 / 64.57 | 80.78 / 67.59 | 86.73 / 75.58 | 88.13 / 77.46 |
enru | 80.39 / 72.45 | 82.6 / 75.19 | 83.09 / 75.77 | 74.54 / 65.96 | 73.71 / 65.9 | 79.43 / 71.83 | 82.54 / 75.36 |
enth | - | - | 76.73 / 71.07 | 72.5 / 67.44 | 67.83 / 62.15 | 75.73 / 71.41 | 77.55 / 73.14 |
entl | - | - | - | 70.4 / 49.57 | 62.96 / 53.84 | 67.74 / 64.78 | 67.41 / 64.79 |
entr | - | - | 76.26 / 73.11 | 66.91 / 63.41 | 63.17 / 58.22 | 75.41 / 72.18 | 77.8 / 75.51 |
enur | - | - | 71.94 / 70.89 | 59.45 / 57.74 | 52.63 / 54.42 | 68.93 / 71.17 | 73.29 / 73.38 |
envi | - | - | 82.48 / 78.01 | 71.57 / 62.53 | 72.89 / 66.38 | 81.39 / 76.66 | 82.03 / 76.85 |
enzh | 76.74 / 71.62 | 79.42 / 74.13 | 79.82 / 73.73 | 71.67 / 65.01 | 73.35 / 64.68 | 79.84 / 73.73 | 80.85 / 75.29 |
Direction | TowerInstruct-7B | TowerInstruct-13B | X-ALMA | Aya-101 | LLaMAX3-Alpaca-8B | GemmaX2-28-2B | GemmaX2-28-9B |
aren | - | - | 43.71 / 87.8 | 35.23 / 85.7 | 37.55 / 86.37 | 45.03 / 87.6 | 49.35 / 88.61 |
aren | - | - | 40.73 / 87.74 | 25.07 / 82.98 | 27.14 / 82.49 | 38.52 / 86.37 | 42.9 / 87.76 |
bnen | - | - | - | 30.06 / 86.45 | 32.52 / 87.65 | 39.29 / 88.74 | 42.76 / 89.71 |
bnen | - | - | - | 20.03 / 82.27 | 23.08 / 81.3 | 32.9 / 86.21 | 34.83 / 86.59 |
csen | - | - | 46.05 / 89.06 | 38.22 / 87.37 | 41 / 87.91 | 45.25 / 88.76 | 47.18 / 89.2 |
csen | - | - | 41.93 / 91.71 | 31.28 / 88.51 | 32.18 / 87.39 | 38.72 / 90.62 | 42.69 / 91.64 |
deen | 49.18 / 89.49 | 50.02 / 89.63 | 49.7 / 89.73 | 42.01 / 88.13 | 44.21 / 88.6 | 49.04 / 89.43 | 49.89 / 89.61 |
deen | 43.68 / 87.82 | 45.4 / 88.16 | 45.95 / 88.45 | 34.74 / 85.17 | 36.33 / 85.07 | 44.85 / 87.91 | 47.13 / 88.5 |
esen | 37.67 / 87.46 | 37.56 / 87.68 | 38.1 / 87.89 | 32.47 / 86.34 | 33.91 / 86.79 | 36.07 / 87.33 | 37.88 / 87.71 |
esen | 33.09 / 86.79 | 33.98 / 87.05 | 33.67 / 87 | 28.2 / 84.96 | 28.58 / 85.13 | 34.34 / 86.81 | 35.45 / 87.23 |
faen | - | - | 41.12 / 88.5 | 33.13 / 86.67 | 35.27 / 87.32 | 41.67 / 88.42 | 44.95 / 89.18 |
faen | - | - | 35.46 / 88.35 | 26.39 / 84.98 | 26.63 / 84.18 | 35.49 / 88 | 38.66 / 88.88 |
fren | 50.98 / 89.55 | 51.32 / 89.68 | 49.89 / 89.56 | 43.78 / 88.27 | 45.2 / 88.47 | 50.62 / 89.48 | 51.18 / 89.64 |
fren | 53.85 / 88.42 | 55.65 / 88.94 | 55.39 / 88.73 | 43.15 / 85.26 | 45 / 85.79 | 54.2 / 88.39 | 57.69 / 89.11 |
heen | - | - | 46.34 / 88.6 | 38.56 / 86.67 | 41.22 / 87.23 | 48.59 / 88.86 | 51.58 / 89.37 |
heen | - | - | 44.45 / 89.26 | 26.74 / 83.65 | 31.36 / 83.35 | 41.96 / 88.48 | 46.26 / 89.26 |
hien | - | - | 44.63 / 89.81 | 35.65 / 87.84 | 37.05 / 88.51 | 45.71 / 89.88 | 49.14 / 90.53 |
hien | - | - | 38.46 / 81.45 | 24.29 / 76.16 | 27.68 / 75.99 | 37.86 / 80.17 | 41.27 / 81.1 |
iden | - | - | 48.7 / 89.85 | 40.92 / 88.37 | 43.07 / 88.83 | 51.21 / 89.95 | 53.27 / 90.37 |
iden | - | - | 49.58 / 92.07 | 38.06 / 89.85 | 38.62 / 89.18 | 51.01 / 91.88 | 52.61 / 92.24 |
iten | 40.54 / 88.4 | 41.17 / 88.46 | 40.61 / 88.46 | 35.28 / 87.12 | 35.77 / 87.39 | 39.84 / 88.24 | 42.06 / 88.59 |
iten | 36.5 / 88.79 | 38.17 / 89.18 | 37.72 / 89.16 | 29.4 / 86.25 | 31.09 / 86.46 | 36.98 / 88.72 | 38.83 / 89.25 |
jaen | - | - | 32.5 / 88.22 | 23.99 / 85.6 | 27.73 / 87.18 | 33.43 / 88.19 | 36.5 / 88.78 |
jaen | - | - | 29.58 / 91.19 | 21 / 88.79 | 22.01 / 88.94 | 30.67 / 90.94 | 33.26 / 91.29 |
kmen | - | - | - | 28.08 / 83.98 | 28.74 / 85.69 | 37 / 87.56 | 41.83 / 88.56 |
kmen | - | - | - | 19.48 / 81.34 | 14.41 / 72.92 | 20.73 / 82.32 | 24.13 / 84.22 |
koen | 34.51 / 88.19 | 36.01 / 88.51 | 34.31 / 88.31 | 27.38 / 86.53 | 29.07 / 87.13 | 35.35 / 88.44 | 38.76 / 89.04 |
koen | 26.39 / 89.4 | 28.39 / 89.87 | 25.74 / 89.27 | 16.34 / 85.94 | 19.62 / 86.79 | 26.7 / 89.24 | 30.37 / 90.13 |
loen | - | - | - | 33.44 / 85.25 | 25.96 / 83.19 | 39.44 / 87.51 | 44.22 / 88.82 |
loen | - | - | - | 28.85 / 83.25 | 10.17 / 62.95 | 27.55 / 82.75 | 31.53 / 84.66 |
msen | - | - | 48.33 / 89.31 | 41.62 / 87.89 | 43.67 / 88.38 | 51.1 / 89.55 | 53.12 / 90.03 |
msen | - | - | 43.23 / 89.9 | 34.21 / 87.11 | 35.7 / 87.39 | 46.09 / 89.82 | 47.03 / 89.89 |
myen | - | - | - | 20.43 / 80.38 | 22.75 / 84.73 | 29.79 / 86.58 | 36.14 / 88.21 |
myen | - | - | - | 15.24 / 84.99 | 9.9 / 73.82 | 15.7 / 85.69 | 20.14 / 88.15 |
nlen | 37.61 / 87.66 | 38.72 / 87.92 | 38.78 / 88.01 | 33.25 / 86.56 | 34.49 / 86.98 | 37.39 / 87.63 | 39.15 / 87.94 |
nlen | 35.24 / 88.31 | 36.41 / 88.66 | 35.97 / 88.59 | 28.72 / 86.09 | 29.51 / 85.66 | 35.01 / 88.24 | 37.54 / 88.82 |
plen | - | - | 37.07 / 86.79 | 29.93 / 84.96 | 32.51 / 85.43 | 36.39 / 86.44 | 38.49 / 86.91 |
plen | - | - | 31.89 / 89.79 | 24.96 / 86.52 | 25.35 / 85.92 | 31.56 / 89.16 | 33.5 / 89.88 |
pten | 55.26 / 89.91 | 56.17 / 90.09 | 53.48 / 89.93 | 47.06 / 88.58 | 49.4 / 88.93 | 54.95 / 89.83 | 57.09 / 90.16 |
pten | 49.19 / 89.42 | 50.76 / 89.73 | 53.42 / 90.1 | 43.34 / 87.69 | 45.28 / 87.88 | 51.22 / 89.67 | 53.17 / 90.04 |
ruen | 40.74 / 86.93 | 42.76 / 87.23 | 40.69 / 87.11 | 33.68 / 85.43 | 36.32 / 86.13 | 40.63 / 86.8 | 43.17 / 87.28 |
ruen | 38.37 / 89.4 | 40.27 / 89.93 | 39.9 / 89.96 | 31.3 / 86.85 | 33.42 / 87.23 | 38.64 / 89.14 | 41.4 / 90.13 |
then | - | - | 34.65 / 88.24 | 27.69 / 86.2 | 30.73 / 87.33 | 37.19 / 88.59 | 40.33 / 89.25 |
then | - | - | 35.87 / 87.47 | 32.45 / 86.12 | 30.79 / 84.72 | 39.45 / 87.91 | 42.53 / 88.7 |
tlen | - | - | - | 37.04 / 84.2 | 43.35 / 86.34 | 52.91 / 88.18 | 55.97 / 88.92 |
tlen | - | - | - | 23.43 / 80.55 | 27.08 / 81.19 | 37.17 / 84.46 | 38.15 / 84.52 |
tren | - | - | 43.63 / 89.73 | 34.67 / 87.76 | 36.62 / 88.31 | 44.43 / 89.72 | 47.97 / 90.47 |
tren | - | - | 37.61 / 89.9 | 28.61 / 86.93 | 24.55 / 85.32 | 39.25 / 89.74 | 42.17 / 90.52 |
uren | - | - | 38.35 / 87.59 | 30.23 / 84.98 | 32.21 / 86.08 | 38.91 / 87.62 | 43.15 / 88.61 |
uren | - | - | 29.56 / 83.38 | 17.78 / 77.56 | 20 / 75.32 | 28.43 / 82.54 | 30.89 / 83.93 |
vien | - | - | 41.5 / 88.03 | 35.02 / 86.09 | 36.71 / 86.95 | 42.73 / 87.98 | 45.45 / 88.53 |
vien | - | - | 44.21 / 89.64 | 30.71 / 84.93 | 36.53 / 86.59 | 44.67 / 89.4 | 46.65 / 89.96 |
zhen | 33.27 / 86.99 | 34.56 / 87.32 | 33.94 / 87.32 | 24.85 / 84.38 | 29.9 / 86.4 | 34.35 / 87.29 | 36.37 / 87.62 |
zhen | 33.9 / 87.77 | 36.35 / 88.45 | 34.43 / 88.04 | 24.24 / 84.81 | 26.96 / 85.4 | 39.39 / 88.88 | 41.6 / 89.18 |
Direction | TowerInstruct-7B | TowerInstruct-13B | Aya-101 | LLaMAX3-Alpaca-8B | GemmaX2-28-2B | GemmaX2-28-9B |
arzh | - | - | 19.62 / 81.35 | 21.34 / 82.33 | 30.33 / 85.23 | 34.11 / 86.37 |
arzh | - | - | 16.31 / 79.75 | 18.35 / 80.54 | 23.2 / 82.56 | 27.62 / 84.39 |
bnzh | - | - | 15.91 / 79.83 | 19.3 / 83.08 | 27.13 / 85.48 | 31.43 / 87.14 |
bnzh | - | - | 11.43 / 76.82 | 14.78 / 77.21 | 20.71 / 81.4 | 24.57 / 82.68 |
cszh | - | - | 19.73 / 82.54 | 22.82 / 83.86 | 31.8 / 86.71 | 35.03 / 87.46 |
cszh | - | - | 17.93 / 85.82 | 18.86 / 86.4 | 23.55 / 87.96 | 28.17 / 89.81 |
dezh | 28.43 / 85.96 | 30.43 / 86.71 | 21.55 / 83.32 | 23.39 / 84.54 | 33.77 / 87.24 | 35.49 / 87.78 |
dezh | 23.93 / 83.48 | 26.53 / 84.66 | 18.52 / 81.28 | 21.08 / 82.65 | 26.56 / 84.19 | 29.29 / 85.39 |
eszh | 25.82 / 85.83 | 27.93 / 86.68 | 24.24 / 84.81 | 26.96 / 85.4 | 30.93 / 87.33 | 32.84 / 87.9 |
eszh | 21.34 / 84.21 | 22.84 / 84.76 | 24.85 / 84.38 | 29.9 / 86.4 | 23.51 / 84.77 | 25.45 / 85.54 |
enzh | 33.9 / 87.77 | 36.35 / 88.45 | 19.96 / 83.76 | 21.3 / 84.33 | 39.39 / 88.87 | 41.64 / 89.19 |
enzh | 33.27 / 86.99 | 34.56 / 87.32 | 17.63 / 82.42 | 19.42 / 83.81 | 34.23 / 87.3 | 36.37 / 87.63 |
fazh | - | - | 18.66 / 82.13 | 20.83 / 83.74 | 29.39 / 86.16 | 33.12 / 87.17 |
fazh | - | - | 15.92 / 81.43 | 17.85 / 82.72 | 23.56 / 85.04 | 26.12 / 86.01 |
frzh | 28.84 / 86.2 | 31.18 / 86.94 | 21.77 / 83.83 | 23.81 / 84.59 | 33.59 / 87.34 | 35.68 / 87.94 |
frzh | 30.07 / 83.92 | 31.55 / 84.96 | 24.29 / 80.98 | 26.03 / 83.06 | 31.93 / 84.77 | 35.25 / 85.63 |
hezh | - | - | 19.48 / 81.98 | 21.91 / 82.77 | 31.13 / 85.78 | 35.14 / 86.95 |
hezh | - | - | 13.92 / 79.22 | 18.36 / 81.24 | 23.12 / 83.65 | 28.53 / 86.03 |
hizh | - | - | 19.77 / 82.8 | 20.59 / 83.04 | 29.74 / 86.05 | 33.27 / 87.18 |
hizh | - | - | 13.86 / 68.98 | 16.89 / 70.61 | 21.54 / 72.62 | 25.41 / 74.49 |
idzh | - | - | 19.99 / 82.24 | 23.17 / 83.98 | 33.4 / 86.96 | 35.79 / 87.7 |
idzh | - | - | 19.72 / 86.1 | 20.77 / 86.26 | 30.08 / 88.61 | 32.34 / 89.21 |
itzh | 26.61 / 86.14 | 29.15 / 86.9 | 19.82 / 83.5 | 22.21 / 84.43 | 31.61 / 87.3 | 33.93 / 87.89 |
itzh | 22.33 / 86.17 | 23.68 / 86.71 | 17.38 / 83.25 | 19.88 / 84.92 | 24.38 / 86.4 | 27.14 / 87.36 |
jazh | - | - | 18.68 / 85.02 | 19.6 / 85.48 | 28.88 / 87.81 | 32.02 / 88.6 |
jazh | - | - | 16.86 / 87.91 | 15.73 / 88.04 | 22.28 / 89.75 | 26.1 / 90.55 |
kmzh | - | - | 7.71 / 69.75 | 18.11 / 81.95 | 26.48 / 85.46 | 30.82 / 86.78 |
kmzh | - | - | 15.84 / 78.6 | 9.15 / 67.7 | 18.61 / 80.24 | 21.99 / 81.93 |
kozh | 24.11 / 85.1 | 27.83 / 86.64 | 19.58 / 83.73 | 21.32 / 84.95 | 29.63 / 86.98 | 32.62 / 88 |
kozh | 17.69 / 86.5 | 20.11 / 87.41 | 11.92 / 83.3 | 14.37 / 85.51 | 19.75 / 87.06 | 22.77 / 88.13 |
lozh | - | - | 17.09 / 80.82 | 13.85 / 79.06 | 26.95 / 85.08 | 31.63 / 86.91 |
lozh | - | - | 18.84 / 80.25 | 6.28 / 58.96 | 20.59 / 80.34 | 25.93 / 82.88 |
mszh | - | - | 18.84 / 81.08 | 22.45 / 82.88 | 32.8 / 86.53 | 35.42 / 87.41 |
mszh | - | - | 17.94 / 82.92 | 18.16 / 83.91 | 26.87 / 86.44 | 28.57 / 86.72 |
myzh | - | - | 9.01 / 75.51 | 12.78 / 80.9 | 21.69 / 83.92 | 27.63 / 86.06 |
myzh | - | - | 11.44 / 81.88 | 7.3 / 67.8 | 11.32 / 81.5 | 15.21 / 84.47 |
nlzh | 25.34 / 84.97 | 27.25 / 85.99 | 18.37 / 81.77 | 20.97 / 83.19 | 29.85 / 86.28 | 31.65 / 86.93 |
nlzh | 22.46 / 85.32 | 23.31 / 85.84 | 17.88 / 82.86 | 19.76 / 84.14 | 23.63 / 85.4 | 26.92 / 86.43 |
plzh | - | - | 19.48 / 82.3 | 20.77 / 83.18 | 29.6 / 85.95 | 32.2 / 86.77 |
plzh | - | - | 16.36 / 84.62 | 17.9 / 85.49 | 23.38 / 87.84 | 25.72 / 89.03 |
ptzh | 29.02 / 86.37 | 31.32 / 87.19 | 21.02 / 83.84 | 23.65 / 84.7 | 34.24 / 87.73 | 36.88 / 88.35 |
ptzh | 26.18 / 85.91 | 28.02 / 86.73 | 22.4 / 83.97 | 25.19 / 85.53 | 30.37 / 86.78 | 33.01 / 87.38 |
ruzh | 26.81 / 84.94 | 29.46 / 85.66 | 20.65 / 82.63 | 22.75 / 83.3 | 31.67 / 86.15 | 33.73 / 86.84 |
ruzh | 22.88 / 86.18 | 26.17 / 87.99 | 19.18 / 85.34 | 22.02 / 86.89 | 25.58 / 87.42 | 28.57 / 88.82 |
thzh | - | - | 19.08 / 83.36 | 21.09 / 85.05 | 29.61 / 87.11 | 32.53 / 88.16 |
thzh | - | - | 25.67 / 83.94 | 24.4 / 82.97 | 31.55 / 85.87 | 35.18 / 86.75 |
tlzh | - | - | 18.78 / 78.78 | 21.36 / 81.07 | 32.3 / 85.05 | 35.63 / 86.02 |
tlzh | - | - | 13.76 / 77.03 | 15.74 / 77.96 | 22.2 / 81 | 24.19 / 81.87 |
trzh | - | - | 19.02 / 81.57 | 21.59 / 83.16 | 30.97 / 86.17 | 34.79 / 87.29 |
trzh | - | - | 16.5 / 81.75 | 13.94 / 81.37 | 23.47 / 84.77 | 26.43 / 86.05 |
urzh | - | - | 16.77 / 79.86 | 19.15 / 81.54 | 26.85 / 84.84 | 31.67 / 86.48 |
urzh | - | - | 10.98 / 73.4 | 13.57 / 73.03 | 17.71 / 77.73 | 21.42 / 79.73 |
vizh | - | - | 20.43 / 84.24 | 22 / 84.89 | 32.15 / 87.43 | 34.26 / 87.95 |
vizh | - | - | 21.31 / 83.98 | 24.81 / 85.98 | 31.67 / 88.35 | 33.79 / 88.8 |